Below are some data from the land of milk and honey. Compute nominal GDP, real GDP,...

  1. Below are some data from the land of milk and honey.

    1. Compute nominal GDP, real GDP, and the GDP deflator for each year, using 2008 as the base year.

    2. Compute the percentage change in nominal GDP, real GDP, and the GDP deflator in 2009 and 2010 from the preceding year. For each year, identify the variable that does not change. Explain in words why your answers make sense.

    3. Did economic well-being rise more in 2009 or 2010? Explain

Year

Price of Milk

Quantity of Milk (litres)

Price of Honey

Quantity of Honey (litres)

2008

$1

100

$2

50

2009

$1

200

$2

100

2010

$2

200

$4

100
Answer #1

a) Nominal GDP

2008: 1*100 + 2*50=$200

2009: 1*200 + 2*100=$400

2010: 2*200 + 4*100=$800

Real GDP

2008: 1*100+2*50=$200

2009: 1*200+2*100=$400

2010: 1*200+2*100=$400

GDP Deflator = Nominal GDP/Real GDP*100

2008:100

2009: (400/400)*100=100

2010: (800/400)*100=200

b) Percentage change in NGDP =

2009: (400-200/200)*100=100%

2010: (800-400/400)*100=100%

Percentage change in RGDP =

2009:(400-200/200)*100=100%

2010: (400-400/400)*100=0%

Percentage change in GDP Deflator

2009: (100-100/100)*100=0%

2010: (200-100/100)*100=100%

The quantity did not change from 2009 to 2010 so the percentage change in real GDP is zero in 2010. Similarly, price did not change from 2008-2009 so the percentage change in GDP Deflator is zero in the year 2009.

c) The economic well being rose more in 2009 because the real GDP increased whereas in 2010, it remained constant as the quantity did not change.